Hi, i have this reaction: 4 NH3 + 3 O2 → 2 N2 + 6 H2O and I only know the moles of NH3 that are: 0.014502748.

how do i get all the others?


1
Expert's answer
2021-07-09T05:15:17-0400

4 NH3 + 3 O→ 2 N2 + 6 H2O

According to the equation:

n (O2) = 3/4 x n(NH3) = 3/4 x 0.014502748 = 0.01087706 mol

n (N2) = 2/4 x n(NH3) = 2/4 x 0.014502748 = 0.00725137 mol

n (H2O) = 6/4 x n(NH3) = 6/4 x 0.014502748 = 0.02175411 mol